OpenStudy (anonymous):

In a large bag of skittles , each of the five (red, orange,yellow,blue,green) occurs with the same probability .you reach in select 2 candies Find the probability that exactly one of the candies is orange ?

OpenStudy (anonymous):

Since the bag has 5 colors (red orange yellow blue and green) it has a denominator of 5 yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And we're saying if we pick two, and only one of them is a certain color

OpenStudy (anonymous):

If we were picking 1 then it would be 1/5 but we're picking 2, how do we account for that?

OpenStudy (anonymous):

since you have 5 different colors and 5 of each color, you have 25 skittles in a bag.

OpenStudy (anonymous):

Is it more or less likely?

OpenStudy (anonymous):

5/25=1/5...

OpenStudy (anonymous):

@anna113 is it more or less likely if we pick 2 instead of 1?

OpenStudy (anonymous):

im not sure i guess more likely ??? @LearningIsAwesome

OpenStudy (anonymous):

im not good in probabilities

OpenStudy (anonymous):

Well we're adding another condition, the condition that we don't pick two of the same color

OpenStudy (anonymous):

is it 2/5 @LearningIsAwesome

OpenStudy (anonymous):

That would mean it's more likely to happen, remember we're not only saying it's a 1/5 chance, but also it we cannot pick it twice

OpenStudy (anonymous):

confused @LearningIsAwesome

OpenStudy (anonymous):

Okay, so the higher your numerator is the higher the probability right? cause 1/5=20% and 2/5=40% yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And here we could be saying we're going to pick just one, and that'd be 5/25=1/5=20% right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

We also could say the probability of just getting that color when picking twice, one or two, and the probability would go up to be 2/5=40% cause we're picking twice right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

but we're not saying those things, we're saying we will pick twice(driving up the probability) but it HAS to be only one(driving down the probability) makes sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Given that each of the five colors have equal probabilty so each has probability P(red)=P(orange)=P(yellow)=P(blue)=P(green)=15 Do you get this part?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

And do you know combinations?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I know part of this but I'm not the best at probability so I wouldn't trust my calculations, I know a few people who can though

OpenStudy (anonymous):

I hope I made it a bit clearer

OpenStudy (anonymous):

yea thanks @LearningIsAwesome

OpenStudy (anonymous):

no problem here I'll tag some people

OpenStudy (anonymous):

@ganeshie8 @phi you guys know this better than I do

OpenStudy (phi):

to get one orange, you picked either Orange + something else or something else + Orange The probability of picking an Orange is 1/5 the probability of picking something else is 4/5 so Pr(O and other)= Pr(O)*Pr(other) = 1/5 * 4/5 = 4/25 the other order other then O also gives 4/25 in total 4/25 + 4/25 = 8/25

OpenStudy (anonymous):

thanks @phi

OpenStudy (phi):

another way to do this is say, Pr( one Orange) = 1 - Pr(2 Oranges) - Pr(no Oranges) Pr (2 Oranges) is 1/5*1/5= 1/25 Pr (no Orange) is 4/5 * 4/5 = 16/15 and Pr( 1 Orange) = 1 - 1/25 - 16/25 = 1 - 17/25 = 25/25 - 17/25 = 8/25